Two seconds after you begin observing a swinging pendulum , the tip of the pendulum is at its rightmost position of 5m from center. Three seconds after that, it is at its leftmost position (5m from center) Determine an equation that models the horizontal position over time.
1
Expert's answer
2019-05-29T09:31:04-0400

Simple Pendulum Equation


"x(t)=A*sin(\omega*t+\phi)"

where

"A=5"

"\omega=2\pi/T=2\pi/5"

"\phi=0"

so


"x(t)=5*sin2\pi/5*t"